The Bellingham Bay Company is planning on increasing its annual dividend by 20 percent a year for the next 2 years and then decreasing the growth rate to 5 percent per year. The company just paid its annual dividend in the amount of $1.00 per share. What is the current value of one share of this stock if the required rate of return is 8 percent?
Answer #1

D1=(1*1.2)=1.2

D2=(1.2*1.2)=1.44

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(1.44*1.05)/(0.08-0.05)

=50.4

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.2/1.08+1.44/1.08^2+50.4/1.08^2

=$45.56(Approx)